# MathGAP: Out-of-Distribution Evaluation on Problems with Arbitrarily Complex Proofs

### v1

@@ 5.3 Nonlinear Depth Generalization
We generate test datasets containing nonlinear problems with comp and comp-eq , having the same form as the problem in Fig. 1. Another example is shown in § B.2. We generate a test set for each depth between $3{-}6$. Thus, the ranged OOD context setting contains examples of depths between $1{-}2$.author=andreas,color=blue!40,size=,fancyline,caption=,]Mention somewhere how depth interacts with width for these problem types
@@ Results.
The results are shown in Fig. 3. All models exhibit a performance trend that tends to zero as depth increases, albeit at different rates. For Mixtral-8x7B, Llama3-8B, and GPT-3.5 Turbo, the performance decreases rapidly across all context modes. These models benefit from in-distribution contexts for low depths, and in some cases, from a range of OOD examples. Llama3-70B and GPT-4o are more robust to this type of OOD distribution shift, [^11] Note that GPT-4o performs better on depth 6 problems than depth 5 problems under the in-distribution prompt. Although we provided the same model id as in the other cases—which according to the documentation points to the original “gpt-4o-2024-05-13” snapshot—this particular test was performed at a later point in time. but for the deepest problems, their performance tends to zero as well. Moreover, it is worth noting that, for many cases, providing in-context examples does not significantly improve performance over zero-shot. On the other hand, for lower depths it is usually beneficial to provide in-distribution contexts. We perform an additional analysis on o1-preview in App. C.
@@ 5.4 Order Generalization
LLMs are known to be sensitive to changes in the order of axioms in logical and mathematical reasoning problems (Chen et al., 2024; Eisape et al., 2024). Here, we present a systematic analysis of their sensitivity to a particular form of order permutation. We consider linear, left-leaning proof trees using comp that have depth 5, and generate problems that deviate from the canonical, left-to-right order of the leaf nodes. In particular, we pick one of the leaf nodes to visit first, and then visit the remaining leaf nodes in left-to-right order. This has the effect of moving one of the sentences in the canonically-ordered problem text to the beginning of the problem. We characterize complexity as the distance of the movement, with the hypothesis that a greater movement from a canonical ordering constitutes a more difficult problem.author=andreas,color=blue!40,size=,fancyline,caption=,]Add a figure to the appendix where we show a linear proof tree and number the canonical orderings and give examples of movements we consider. Similar style as Fig. 1 but less stuff in it. We create five test sets with movement distances between $1{-}5$. We do not include a primitive OOD in-context distribution, the notion of a primitive problem is ill-defined in this case. The range prompt includes movements of all distances between $1{-}5$ except the one present in the test set.
Figure 4: Answer accuracies for generalization to permutations, across models and in-context distributions. We measure complexity as the distance of the movement of a sentences moved to the beginning of the problem, as compared to its position in the canonical ordering of the problem.
@@ Results.
The results are shown in Fig. 4. First, note that all models show lower performance on perturbed problems as compared to non-perturbed ones (compare to Fig. 2). Second, there is a nonlinear relationship between accuracy and distance of the movement, in contrast to the monotonically decreasing relation we expected. More specifically, the performance is the highest for the problems where the sentence is moved from near the beginning or the end of the problem. [^12] This is consistent with findings on retrieval-augmented generation, for which models are better at using information that occurs near the beginning or end of the prompt (Liu et al., 2024a). In addition, including in-context examples gives a larger boost in performance for short and long movements, as compared to medium-length ones.
@@ 6 Conclusion
In this paper, we introduced MathGAP, a framework for evaluating LLMs on math word problems of arbitrary arithmetic proof complexity. MathGAP can flexibly generate synthetic arithmetic word problems with controllable proof tree characteristics. Importantly, this enables studies of OOD generalization that avoid issues of data contamination, since a train-test split where the distributions are different can be created programmatically. In particular, the test set can be made arbitrarily more complex than the training set. We apply MathGAP to study whether LLMs can learn from simple examples in context to solve more complex problems in a test set. All LLMs tested show a decrease in performance as proof complexity increases through depth and/or width. We also find that LLMs are sensitive to the order in which sentences are presented in a particular manner: A problem is harder if a sentence is moved from the middle of the problem, as compared to the beginning or end. We find no clear relationship between the distribution of in-context examples and performance; zero-shot or OOD contexts sometimes yield performance higher than or comparable to in-distribution contexts. Finally, we demonstrate that we can use MathGAP to construct problems that are challenging even for state-of-the-art reasoning models like OpenAI o1—see App. C.
Limitations
Although we make sure to sample from a wide range of agent names, entities, attributes and units, and include multiple templates for the predicates, future work may investigate how our findings generalize to distributions of problems with higher linguistic diversity. It is also beyond the scope of the present study to consider problem texts in languages other than English; however, it is straightforward to extend MathGAP to other languages. In addition, we cannot guarantee that the distributions of our generated problems are different from those of any potential internal data-generating mechanisms from Meta or OpenAI. However, we deem it to be unlikely that they use a similar formalism as we do here, and it gets increasingly unlikely that problems are similar as complexity increases.
While the more complex nonlinear problems are indeed challenging even for the most capable models, we note that the main purpose of our study was not to create a challenging evaluation set for state-of-the-art LLMs. If that is the goal, one could create harder problems by combining several inference rules (e.g., including the rate concept; Opedal et al., 2023), designing a wider variety of proof tree topologies, and presenting the axioms in arbitrary orderings. It is also possible to create more syntactically complex templates and use numbers on which it has been shown to be more challenging for models to perform arithmetic (Razeghi et al., 2022).
@@ Acknowledgments
We thank Alessandro Stolfo, Shehzaad Dhuliawala, Yanick Zengaffinen, and Yilmazcan Ozyurt for useful feedback and discussion. Andreas Opedal acknowledges funding from the Max Planck ETH Center for Learning Systems.
Appendix A Example of a World Model
@@
Appendix B Example Problems
author=yilmazcan,color=white!40,size=,fancyline,caption=,]For each subsection, maybe a small comment about why this problem is depth 6 or 3, or width 6 etc?
@@ Appendix B Example Problems
B.1 Example linear comparison problem (depth 6)
@@
B.2 Example nonlinear comparison problem (depth 3)
@@
B.3 Example partwhole width problem (width 7)
B.4 Permuted problems
@@
Appendix C Additional Analysis on OpenAI o1
@@
Fig. 6 shows additional results on the OpenAI o1-preview model, “o1-preview-2024-09-12”, evaluated on the nonlinear test sets. We only used zero-shot prompting due to the high costs of performing inference with o1-preview. Using a simple, straightforward prompt is also consistent with OpenAI’s current recommendations. We see that the o1 model shows superior performance to all models from the main text. However, as with the other models, the performance decreases as complexity increases.
@@
Importantly, we note that performance can be improved by increasing the limit on the number of output tokens. The results presented in Fig. 6 used a token limit of 4,096, like the experiments in the main paper. We compared the effect of the token limit on an additional test set of 400 nonlinear problems with depth 7. With a token limit of 4,096, the model could only answer $0.25\%$ of the problems correctly. However, with a token limit of 10,000, it answered $76.5\%$ of the problems correctly. These results are noteworthy, and suggests that the new inference technique used by the o1 model is highly effective in decomposing complex problems into smaller steps.
To test the model’s limits, we performed a final evaluation on a subset of 100 of the depth 7 problems where the sentences in each problem were randomly ordered. Recall from § 3.2 that permuting the sentences in random order is valid since all the inference rules used in these problems are commutative. In this experiment, we allowed 25,000 output tokens, which is the current recommendation by OpenAI. On this set, the answer accuracy was only $5\%$. Thus, we conclude that MathGAP is future-proof in the sense that it can generate problems on which current state-of-the-art models fail. We stress, however, that the aim of this paper is not to construct a challenging test set for state-of-the-art models; there are several other ways that MathGAP can be used to create problems that might be even more difficult.
@@
Figure 6: Answer accuracies for generalization to increasing depth and width for nonlinear problems for OpenAI o1-preview under zero-shot prompting.

### v3

@@ 5.3 Nonlinear Depth Generalization
We generate test datasets containing nonlinear problems with comp and comp-eq which have a similar form as the problem in Fig. 1. The procedure is explained in App. D and a more complex, depth $3$ example is given in § E.3. We generate a test set for each depth between $3{-}6$. Thus, the ranged OOD context setting contains examples of depths between $1{-}2$.
@@ Results.
The results are shown in Fig. 3. All models exhibit a performance trend that tends to zero as depth increases, albeit at different rates. For Mixtral-8x7B, Llama3-8B, and GPT-3.5 Turbo, the performance decreases rapidly across all context modes. These models benefit from in-distribution contexts for low depths, and in some cases, from a range of OOD examples. Llama3-70B and GPT-4o are more robust to this type of OOD distribution shift, [^11] Note that GPT-4o performs better on depth 6 problems than depth 5 problems under the in-distribution prompt. Although we provided the same model id as in the other cases—which according to the documentation points to the original “gpt-4o-2024-05-13” snapshot—this particular test was performed at a later point in time. but for the deepest problems their performance tends to zero as well. Thus, the models are able to generalize to some extent but the error rate increases with complexity, suggesting that LLMs are noisy reasoners (Prabhakar et al., 2024). Moreover, it is worth noting that, for many cases, providing in-context examples does not significantly improve performance over zero-shot. On the other hand, for lower depths, it is usually beneficial to provide in-distribution contexts. We perform an additional analysis on o1-preview and DeepSeek-R1 in App. F.
Comparison to linear problems.
Nonlinear problems are more difficult than the linear ones, even when controlling for the number of axioms given in the problem (i.e., its width). To see this, note that a nonlinear problem with depth 3 has 10 axioms (see App. D). This is the same number of axioms as the linear comparison and transfer problems (§ 5.1) with depth 9 and the linear part-whole problems (§ 5.2) with width 10. Comparing the results across those problem sets, we observe considerably lower accuracies for the nonlinear problems in most cases. We believe there are two main explanations for this: (i) Solving nonlinear problems requires keeping intermediate conclusions in memory while proving other intermediate conclusions, before being able to use them for further proof steps. It is thus less predictable where the relevant tokens will be located within the context window as compared to linear problems, in which intermediate conclusions are used immediately in a new proof step. (ii) Nonlinear problems use the comp-eq rule, which is most likely less frequent in the training set. Inspecting the model outputs we find support for both of these explanations—errors show up both when a previously deduced conclusion is needed as well as when a comp-eq is required in a proof step.
@@ 5.4 Order Generalization
LLMs are known to be sensitive to changes in the order of axioms in logical and mathematical reasoning problems (Chen et al., 2024; Eisape et al., 2024). Here, we present a systematic analysis of their sensitivity to a particular form of order permutation. We consider linear, left-leaning proof trees using comp that have depth 5, and generate problems that deviate from the canonical, left-to-right order of the leaf nodes. In particular, we pick one of the leaf nodes to visit first, and then visit the remaining leaf nodes in left-to-right order. This has the effect of moving one of the sentences in the canonically-ordered problem text to the beginning of the problem. An illustration is shown in Fig. 7 and § E.4 gives further examples. We characterize complexity as the distance of the movement, with the hypothesis that a greater movement from a canonical ordering constitutes a more difficult problem. We create five test sets with movement distances between $1{-}5$. We do not include a primitive OOD in-context distribution since the notion of a primitive problem is ill-defined in this case. The range prompt includes movements of all distances between $1{-}5$ except the one present in the test set.
Figure 4: Answer accuracies for generalization to permutations across models and in-context distributions. We measure complexity as the distance of the movement of a sentence to the beginning of the problem, as compared to its position in the canonical ordering of the problem.
@@ Results.
The results are shown in Fig. 4. First, note that all models show lower performance on perturbed problems as compared to non-perturbed ones (compare to Fig. 2), suggesting that they are sensitive to simple perturbations of axiom orderings. Second, there is a nonlinear relationship between accuracy and distance of the movement, in contrast to the monotonically decreasing relation we expected. More specifically, the performance is the highest for the problems where the sentence is moved from near the beginning or the end of the problem. [^12] This is consistent with findings on retrieval-augmented generation, for which models are better at using information that occurs near the beginning or end of the prompt (Liu et al., 2024a). In addition, including in-context examples gives a larger boost in performance for short and long movements, as compared to medium-length ones. Inspecting the model outputs, we observe that the models often make logical mistakes when the sentence that has been moved is needed for the subsequent proof steps. These kinds of mistakes appear to be more common than arithmetic errors.
@@ 6 Conclusion
This paper introduced MathGAP, a framework for evaluating LLMs on math word problems with proofs that are arbitrarily complex. MathGAP can flexibly generate synthetic arithmetic word problems with controllable proof tree characteristics. Importantly, this enables studies of easy-to-hard OOD generalization that avoid issues of data contamination, since a train-test split where the distributions are different can be created programmatically. In particular, the test set can be made arbitrarily more complex than the training set. We applied MathGAP to study whether LLMs can learn from simple examples in context to solve more complex problems in a test set. All models showed a decrease in performance as complexity of proof trees increases through depth and/or width. The LLMs struggled more with wider problems than with deeper problems, and were worse at solving problems with nonlinear shapes as compared to linear ones. We also found that LLMs are sensitive to the order in which sentences are presented. We further demonstrated that we can use MathGAP to construct problems that are challenging even for state-of-the-art reasoning models like OpenAI o1 and DeepSeek-R1—see App. F.
Code to generate problems with MathGAP, including problem types beyond those considered in this paper, can be found in our public code repository. Finally, see App. G for a discussion on limitations of the present work.
@@ Acknowledgments
We thank Alessandro Stolfo, Ryan Cotterell, Shehzaad Dhuliawala, Yanick Zengaffinen, and Yilmazcan Ozyurt for useful discussion. Andreas Opedal acknowledges funding from the Max Planck ETH Center for Learning Systems.
Appendix A Intuition on Logical Forms
A cont logical form expresses the fact that an agent owns a particular entity in some quantity (e.g., “Alice has 5 apples”). A comp logical form expresses how much more some agent owns of a particular entity than some other agent (e.g., “Bob has 3 fewer apples than Alice”). A transfer logical form expresses the transfer of a certain quantity of an entity from one agent to another (e.g., “Alice gave Bob 3 apples”). A partwhole logical form expresses a superset relation between the combined entities of several agents (e.g., Alice and Bob combine their fruits). Finally, a comp-eq logical form expresses that the quantity of a particular comp between two agents is equal to the difference of the quantities of the entity of two other agents (e.g., “The number of apples that Alice has more than Bob is the same as the difference between the number of apples that Charlie and David have”).
@@ Appendix A Intuition on Logical Forms
Appendix B Defining Inference Rules with Natural Deduction
[…]
  1. 1.
    $\exists v_{1}(\texttt{Owner}(v_{1},\texttt{a})\land\texttt{Measure}(v_{1},\texttt{q}_{1})\land\forall x\in v_{1}.\texttt{e}(x))$ by Axiom (1st premise, definition of container as in Opedal et al. 2023, App. D.2.1).
  2. 2.
    $\texttt{Owner}(v_{1},\texttt{a})\land\texttt{Measure}(v_{1},\texttt{q}_{1})\land\forall x\in v_{1}.\texttt{e}(x)$ by $\exists$-elimination [1].
  3. 3.
    $\exists e_{1}(\texttt{ComparisonAdd}(e_{1})\land\texttt{Source}(e_{1},v_{1})\land\texttt{Target}(e_{1},v_{2})\land\texttt{Time}(v_{1})=\texttt{Time}(v_{2})\land\exists r(\texttt{Arg}(e_{1},r)\land\texttt{Measure}(r,\texttt{q}_{2})\land\forall x\in r.\texttt{e}(x)))$ by Axiom (2nd premise, definition of comparison as in Opedal et al. 2023, App. D.2.2).
  4. 4.
    $\texttt{ComparisonAdd}(e_{1})\land\texttt{Source}(e_{1},v_{1})\land\texttt{Target}(e_{1},v_{2})\land\texttt{Time}(v_{1})=\texttt{Time}(v_{2})\land\exists r(\texttt{Arg}(e_{1},r)\land\texttt{Measure}(r,\texttt{q}_{2})\land\forall x\in r.\texttt{e}(x))$ by $\exists$-elimination [3].
  5. 5.
    $\exists r(\texttt{Arg}(e_{1},r)\land\texttt{Measure}(r,\texttt{q}_{2})\land\forall x\in r.\texttt{e}(x))$ by $\land$-elimination [4].
  6. 6.
    $\texttt{Arg}(e_{1},r)\land\texttt{Measure}(r,\texttt{q}_{2})\land\forall x\in r.\texttt{e}(x)$ by $\exists$-elimination [5].
  7. 7.
    $\texttt{Arg}(e_{1},r)\land\forall x\in r.\texttt{e}(x)$ by $\land$-elimination [6].
  8. 8.
    $\texttt{ComparisonAdd}(e_{1})\land\texttt{Arg}(e_{1},r)\land\forall x\in r.\texttt{e}(x)$ by $\land$-introduction [7].
  9. 9.
    $\forall e\forall r(\texttt{ComparisonAdd}(e)\land\texttt{Arg}(e,r)\land\forall x\in r.\texttt{e}(x)\to\exists v\exists o\exists q(\texttt{Target}(e,v)\land\texttt{Owner}(v,o)\land\texttt{Measure}(v,q)\land\forall x\in v.\texttt{e}(x))$ by Axiom (theorem on the existence of containers from existing relations as in Opedal et al. 2023, App. D.4).
  10. 10.
    $\texttt{ComparisonAdd}(e_{1})\land\texttt{Arg}(e_{1},r)\land\forall x\in r.\texttt{e}(x)\to\exists v\exists o\exists q(\texttt{Target}(e_{1},v)\land\texttt{Owner}(v,o)\land\texttt{Measure}(v,q)\land\forall x\in v.\texttt{e}(x))$ by $\forall$-elimination [9].
  11. 11.
    $\exists v\exists o\exists q(\texttt{Target}(e_{1},v)\land\texttt{Owner}(v,o)\land\texttt{Measure}(v,q)\land\forall x\in v.\texttt{e}(x))$ by $\to$-elimination [8, 10].
  12. 12.
    $\texttt{Target}(e_{1},v_{2})\land\texttt{Owner}(v_{2},\texttt{b})\land\texttt{Measure}(v_{2},\texttt{q}_{t})\land\forall x\in v_{2}.\texttt{e}(x)$ by $\exists$-elimination [11].
  13. 13.
    $\texttt{Owner}(v_{2},\texttt{b})\land\texttt{Measure}(v_{2},\texttt{q}_{t})\land\forall x\in v_{2}.\texttt{e}(x)$ by $\land$-elimination [12].
  14. 14.
    $\forall e\forall v_{s}\forall v_{t}\forall m_{s}\forall m_{t}\forall r(\texttt{ComparisonAdd}(e)\land\texttt{Arg}(e,r)\land\texttt{Source}(e,v_{s})\land\texttt{Target}(e,v_{t})\land\texttt{Measure}(v_{s},m_{s})\land\texttt{Measure}(v_{t},m_{t})\to m_{s}+r=m_{t})$ by Axiom (theorem defining the semantics of comparison as in Opedal et al. 2023, App. D.3).
  15. 15.
    $\texttt{ComparisonAdd}(e_{1})\land\texttt{Arg}(e_{1},r)\land\texttt{Source}(e_{1},v_{1})\land\texttt{Target}(e_{1},v_{2})\land\texttt{Measure}(v_{1},\texttt{q}_{1})\land\texttt{Measure}(v_{2},\texttt{q}_{t})\to\texttt{q}_{1}+\texttt{q}_{2}=\texttt{q}_{t}$ by $\forall$-elimination [14].
  16. 16.
    $\texttt{Measure}(v_{1},\texttt{q}_{1})$ by $\land$-elimination [2].
  17. 17.
    $\texttt{Measure}(v_{2},\texttt{q}_{t})$ by $\land$-elimination [13].
  18. 18.
    $\texttt{ComparisonAdd}(e_{1})\land\texttt{Arg}(e_{1},r)$ by $\land$-elimination [8].
  19. 19.
    $\texttt{Source}(e_{1},v_{1})\land\texttt{Target}(e_{1},v_{2})$ by $\land$-elimination [4].
  20. 20.
    $\texttt{ComparisonAdd}(e_{1})\land\texttt{Arg}(e_{1},r)\land\texttt{Source}(e_{1},v_{1})\land\texttt{Target}(e_{1},v_{2})\land\texttt{Measure}(v_{1},\texttt{q}_{1})\land\texttt{Measure}(v_{2},\texttt{q}_{t})$ by $\land$-introduction [4, 16, 17, 18, 19].
  21. 21.
    $\texttt{q}_{1}+\texttt{q}_{2}=\texttt{q}_{t}$ by $\to$-elimination [15, 20].
  22. 22.
    $\texttt{Owner}(v_{2},\texttt{b})\land\texttt{Measure}(v_{2},\texttt{q}_{1}+\texttt{q}_{2})\land\forall x\in v_{2}.\texttt{e}(x)$ by $=$-elimination [13, 21].
  23. 23.
    $\exists v_{2}(\texttt{Owner}(v_{2},\texttt{b})\land\texttt{Measure}(v_{2},\texttt{q}_{1}+\texttt{q}_{2})\land\forall x\in v_{2}.\texttt{e}(x))$ by $\exists$-introduction [22].
Note that the conclusion (step 22) is the first-order logic equivalent of cont(b, q1 + q2, e), which is the conclusion of the comparison inference rule.
Appendix C More Details on Experiments
Generated problems.
The quantities are sampled uniformly at random from the range $2{-}20$. We use a handwritten list of 52 English-language names for most problems, with the exception of the larger nonlinear problems for which we use a larger list of 4,945 predominantly English-language names from https://github.com/dominictarr/random-name. For the entities we use a handwritten word list with 51 items. We restrict each problem to have one entity, We determine uniformly at random whether the problem has attributes, units, or neither of them.
Prompts.
[…]
Models.
The maximum context windows for Mixtral-8x7B, Llama3 8B, Llama3 70B, GPT-3.5 Turbo and GPT-4o are 65k, 8k, 8k, 16k, 128k, respectively. Combined with a maximum number of output tokens of 4,096, the models have a theoretically sufficient context size to solve all test problems. The largest test problems (nonlinear depth 6) have around 2.5k tokens, including CoT annotation.
For GPT-3.5 Turbo and GPT-4o we used “gpt-3.5-turbo-0125” and “gpt-4o-2024-05-13” as the model ids for all experiments.
Appendix D Nonlinear Comparison Problems
In this section we explain why we have introduced the comp-eq predicate to generate nonlinear trees. We also give the procedure for generating the nonlinear problems used for the experiments in § 5.3.
Motivating example.
[…]
Now, say we are following the top-down generation procedure described in § 4.1 and wish to expand the tree to higher depths. Note that the two leaf nodes labeled with comp are impossible to expand further. For instance, the comp(B, A, 3, a) would be deduced by the two premises cont(A, 5, a) and cont(B, 8, a), but those two logical forms are already present elsewhere in the tree and are thus not available. This leaves only the two cont nodes for further expansion. However, any further subtrees of the cont nodes will be linear, since none of the comp nodes generated along the way will be able to be expanded. [^13] This reasoning assumes that we only use the inference rules with cont and comp from § 3.1. Deeper nonlinear problems could be achieved by using the partwhole rule as well.
[…]
Procedure.
We want to sample a nonlinear proof tree with depth $D$. We start by sampling either a cont node or a comp node uniformly at random to determine the root of the tree. Next, we sample an inference rule for which the root logical form is the conclusion. We then sample inference rules from the child nodes in breadth-first order until all leaves except the ones labeled with comp-eq (which cannot be expanded) have depth $D$. For comp nodes, there is only one inference rule possible (the third row in § 3.1). For cont nodes we sample a comp rule (i.e., the first row in § 3.1) if the cont node has depth $D-1$, and a comp-eq rule (i.e., the fifth row in § 3.1) otherwise. This ensures that we can continue to grow the tree nonlinearly until the specified depth, following the reasoning given in the paragraph above.
Note that the width of the proof trees generated with this procedure will be the number of leaf nodes in a perfect binary tree with depth $D$, i.e., $2^{D}$, plus the number of nodes labeled with comp-eq . Thus, a depth 2 problem will have a width of either 4 or 5, a depth 3 problem will have a width of 10, and a depth 4 problem will have a width of either 20 or 21. In general, the width $W(D)$ for a proof tree of depth $D$ is:
$W(D)=2^{D}+\sum_{d=0}^{D-2}f_{c}(d),$
where $f_{c}(d)$ gives the number of comp-eq nodes as a function of the level $d$, following the recursion
$f_{c}(0)=\begin{cases}1\quad\text{if root is {cont} missing}\\ 0\quad\text{if root is {comp} missing}\end{cases}$
$f_{c}(1)=\begin{cases}1\quad\text{if root is {cont} missing}\\ 2\quad\text{if root is {comp} missing}\end{cases}$
$f_{c}(d)=f_{c}(d-1)+2f_{c}(d-2),d\geq 2.$
Appendix E Example Problems and Illustrations
Figure 6: Example of a linear problem using comp with depth 3 and width 4. This problem is linear because each parent node in the tree has at most one child that is not a leaf node. Compare to the nonlinear problem shown in Fig. 1.
@@ Appendix E Example Problems and Illustrations
E.1 Example linear comparison problem (depth 6)
@@
E.2 Example partwhole width problem (width 7)
E.3 Example nonlinear comparison problem (depth 3)
@@
E.4 Permuted problems
Figure 7: The same proof tree as Fig. 6 but with a permuted ordering as in the experiments discussed in § 5.4. In this specific example, the leaf node labeled with comp(L, J, 11, apple) is visited first (the other nodes follow post-order traversal). This translates into moving the sentence “Lucy has 11 more apples than John” to the front of the problem. As compared to the canonically ordered problem shown in Fig. 6, the sentence has moved two steps.
@@ E.4 Permuted problems
Appendix F Additional Analysis on Reasoning Models
@@
Fig. 8 shows additional results on the OpenAI o1-preview model (“o1-preview-2024-09-12”) and DeepSeek-R1 (DeepSeek-AI, 2025), evaluated on the nonlinear test sets. We only used zero-shot prompting due to the high costs of performing inference on these models. Using a simple, straightforward prompt is also consistent with OpenAI’s current recommendations. We see that these models show superior performance to all models from the main text, with o1 generalizing better than R1. However, as with the other models, the performance decreases as complexity increases.
@@
Importantly, we noted that performance can be drastically improved by increasing the limit on the number of output tokens. The results presented in Fig. 8 used a token limit of 4,096, like the experiments in the main paper. For o1, we compared the effect of the token limit on an additional test set of 400 nonlinear problems with depth 7. With a token limit of 4,096, the model could only answer $0.25\%$ of the problems correctly. However, with a token limit of 10,000, it answered $76.5\%$ of the problems correctly. For R1, the performance on problems with depth 6 is increased to $77.5\%$ when using a token limit of 10,000. These results are noteworthy and suggests that the inference techniques used by these models is highly effective for decomposing complex problems into smaller steps.
@@
Figure 8: Answer accuracies for generalization to increasing depth and width for nonlinear problems for OpenAI o1-preview and DeepSeek-R1 under zero-shot prompting and a maximum context length of 4,096 tokens.
To test the limits of these models, we performed a final evaluation on a subset of 100 of the depth 7 problems where the sentences in each problem were randomly ordered (apart from the question). Recall from § 3.2 that permuting the sentences in random order is valid since all the inference rules used in our nonlinear problems are commutative. In this experiment, we allowed 25,000 output tokens for both models, which is the current recommendation by OpenAI. On this set, the answer accuracy was only $5.0\%$ for o1 and $11.0\%$ for R1. Thus, we conclude that MathGAP is future-proof in the sense that it can generate problems on which current state-of-the-art reasoning models fail. We stress, however, that the aim of this paper is not to construct a challenging test set for state-of-the-art models; there are several other ways that MathGAP can be used to create problems that might be even more difficult.
Appendix G Limitations
We did not prioritize linguistic diversity in our problem sets; future work may investigate how our findings generalize to distributions of problems with higher linguistic diversity. We did also not consider or study the possibility of token biases (Jiang et al., 2024b). It is beyond the scope of the present study to consider problem texts in languages other than English; however, it is straightforward to extend MathGAP to other languages. In addition, we cannot guarantee that the distributions of our generated problems are different from those of any potential internal data-generating mechanisms from Meta or OpenAI. However, we deem it to be unlikely that they use a similar formalism as we do here, and it gets increasingly unlikely that problems are similar as complexity increases.
While the more complex nonlinear problems are indeed challenging even for the most capable models, we note the purpose of our study was not to create a challenging evaluation set for state-of-the-art LLMs. If that is the goal, one could create harder problems by combining several inference rules (e.g., additionally including the rate concept; Opedal et al., 2023), allowing a wider variety of proof-tree shapes, and presenting the axioms in arbitrary orderings. It is also possible to create more syntactically complex templates (and paraphrase those using LLMs) and use numbers on which it is more challenging for models to perform arithmetic (Razeghi et al., 2022).